Solve the exact differential equation = ( x - 2ey ) dy + y + ( xsin(x) ) dx = 0


1
Expert's answer
2022-02-07T16:34:31-0500
"P(x,y)=y+x\sin x, \dfrac{\partial P}{\partial y}=1"

"Q(x,y)=x-2e^y, \dfrac{\partial Q}{\partial x}=1"

"\dfrac{\partial P}{\partial y}=1=\dfrac{\partial Q}{\partial x}"

Write the system of two differential equations that define the function "u(x,y)"


"\dfrac{\partial u}{\partial x}=y+x\sin x"

"\dfrac{\partial u}{\partial y}=x-2e^y"

"u=\int(x-2e^y)dy+\varphi(x)"

"u=xy-2e^y+\varphi(x)"


"\dfrac{\partial u}{\partial x}=y+\varphi'(x)=y+x\sin x"

"\varphi'(x)=x\sin x"


"\varphi(x)=\int x\sin xdx"

"\varphi(x)=-x\cos x+\sin x+C"


"u=xy-2e^y-x\cos x+\sin x+C"

"-xy+2e^y+x\cos x-\sin x=C"